# references.bib

### v1

@@ Appendix A Full List of Wheat and Barley Papers by Weed Species
Stellaria media Common Chickweed 35 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, zhuang2022evaluation, alirezazadeh2023comparative, anderegg2023farm, bharathi2023plant, dheeraj2023using, guo2023identification, liu2023semi, mishra2023comparative, mu2023densenet, rahman2023new]
Taraxacum officinale Common Dandelion 1 [anderegg2023farm]
Trifolium pratense Red Clover 2 [dyrmann2017roboweedsupport, dyrmann2018using]
@@
Tripleurospermum inodorum Scentless Mayweed 32 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new, thomas2023weakly]
Urtica urens Small Nettle 2 [dyrmann2017roboweedsupport, dyrmann2018using]
@@
Veronica hederifolia Ivy-leaved Speedwell 2 [pflanz2018weed, anderegg2023farm]
@@
Veronica persica Common field-speedwell 5 [sukumar2016weed, dyrmann2017roboweedsupport, dyrmann2018using, anderegg2023farm, liu2023semi]
@@
Viola arvensis Field Pansy 4 [dyrmann2017roboweedsupport, dyrmann2018using, pflanz2018weed, anderegg2023farm]
Vulpia myuros Rattail Fescue 1 [jensen2021predictive]
Xanthium strumarium 2 [wang2018low, wang2020transfer]
[…]

### v2

@@ Appendix A Full List of Wheat and Barley Papers by Weed Species
Stellaria media Common Chickweed 35 dyrmann2015roboweedsupport ; dyrmann2017roboweedsupport ; chavan2018agroavnet ; dyrmann2018using ; binguitcha2019crops ; elnemr2019convolutional ; jiang2019novel ; espejo2020improving ; fu2020field ; rahman2020performance ; skacev2020development ; trong2020late ; espejo2021combining ; espejo2021testing ; farkhani2021weed ; mounashree2021automated ; siddiqui2021neural ; trong2021yielding ; damilare2022weed ; kumar2022weed ; kundur2022deep ; mu2022faster ; reddy2022weed ; srivastava2022role ; tara2022machine ; zhuang2022evaluation ; alirezazadeh2023comparative ; anderegg2023farm ; bharathi2023plant ; dheeraj2023using ; guo2023identification ; liu2023semi ; mishra2023comparative ; mu2023densenet ; rahman2023new
Taraxacum officinale Common Dandelion 1 anderegg2023farm
Trifolium pratense Red Clover 2 dyrmann2017roboweedsupport ; dyrmann2018using
@@
Tripleurospermum inodorum Scentless Mayweed 32 dyrmann2015roboweedsupport ; dyrmann2017roboweedsupport ; chavan2018agroavnet ; dyrmann2018using ; binguitcha2019crops ; elnemr2019convolutional ; jiang2019novel ; espejo2020improving ; fu2020field ; rahman2020performance ; skacev2020development ; trong2020late ; espejo2021combining ; espejo2021testing ; farkhani2021weed ; mounashree2021automated ; siddiqui2021neural ; trong2021yielding ; damilare2022weed ; kumar2022weed ; kundur2022deep ; mu2022faster ; reddy2022weed ; srivastava2022role ; tara2022machine ; bharathi2023plant ; dheeraj2023using ; guo2023identification ; mishra2023comparative ; mu2023densenet ; rahman2023new ; thomas2023weakly
Urtica urens Small Nettle 2 dyrmann2017roboweedsupport ; dyrmann2018using
@@
Veronica hederifolia Ivy-leaved Speedwell 2 pflanz2018weed ; anderegg2023farm
@@
Veronica persica Common field-speedwell 5 sukumar2016weed ; dyrmann2017roboweedsupport ; dyrmann2018using ; anderegg2023farm ; liu2023semi
@@
Viola arvensis Field Pansy 4 dyrmann2017roboweedsupport ; dyrmann2018using ; pflanz2018weed ; anderegg2023farm
Vulpia myuros Rattail Fescue 1 jensen2021predictive
Xanthium strumarium 2 wang2018low ; wang2020transfer